# A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare

### v1

@@ Appendix B Comparison of Models
SOLAR (10.7B) 2023 Upstage AI General-purpose NLP, text generation, and fine-tuning for specific domains. Explainability
@@
InternLM2 (7B) 2023 Shanghai AI Laboratory (China) General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
@@
Llama3-( 8B and 70B) 2024 Meta General-purpose NLP, text generation, and fine-tuning for specific applications. Privacy
@@
CodeLlama-( 7B, 13B, and 34B) 2023 Meta Code generation, code completion, and programming assistance. Privacy
@@
Mixtral-8x7B and 8x22B 2023 Mistral AI General-purpose NLP, text generation, and fine-tuning for specific domains. Privacy
@@
Qwen-(7B, 14B, 32B, 72B)-Chat 2023 Alibaba Chat-oriented NLP, conversational AI, and text generation. Privacy
@@
GLM-4 2024 Tsinghua University Advanced NLP, text generation, and multi-modal tasks. Privacy
@@
Table 2: Detailed Comparison of GPT Models Evaluated for Trust in Healthcare LLMs, Including Model Name, Release Year, Institution, Primary Tasks (e.g., Medical Question-Answering, Clinical Decision Support, Biomedical Text Summarization, Medical Report Generation), and Key Trustworthiness Dimensions (Truthfulness, Privacy, Safety, Robustness, Fairness and Bias, Explainability).

### v2

@@ Appendix C Comparison of Models
SOLAR (10.7B) 2023 Upstage AI Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific domains. Explainability
@@
InternLM2 (7B) 2023 Shanghai AI Laboratory (China) Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
@@
Llama3-( 8B and 70B) 2024 Meta Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific applications. Privacy, Explainability
@@
CodeLlama-( 7B, 13B, and 34B) 2023 Meta Open-source Decoder-only Code generation, code completion, and programming assistance. Privacy
@@
Mixtral-8x7B and 8x22B 2023 Mistral AI Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific domains. Privacy
@@
Qwen-(7B, 14B, 32B, 72B)-Chat 2023 Alibaba Open-source Decoder-only Chat-oriented NLP, conversational AI, and text generation. Privacy
@@
GLM-4 2024 Tsinghua University Open-source Encoder-decoder Advanced NLP, text generation, and multi-modal tasks. Privacy
@@
Table 2: Detailed Comparison of GPT Models Evaluated for Trust in Healthcare LLMs, Including Model Name, Release Year, Institution, Openness, Architecture, Primary Tasks (e.g., Medical Question-Answering, Clinical Decision Support, Biomedical Text Summarization, Medical Report Generation), and Key Trustworthiness Dimensions (Truthfulness, Privacy, Safety, Robustness, Fairness and Bias, Explainability).